Multiply and reduce to lowest form (if possible):
(2/7) × (7/9)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (2 × 7)/ (7 × 9)
= (2 × 1)/ (1 × 9)
= (2/9)

Thanks